A 46.0-kg woman wearing high-heeled shoes is invited into a home in which the kitchen has vinyl floor covering. The heel on each shoe is circular and has a radius of 0.540 cm.

a)) If the woman balances on one heel, what pressure does she exert on the floor? Check the definition of pressure N/m2 please

Explain your answer.

thank you!
Answer #1

Pressure = Force ÷ area
Force = weight = m * g = 46 * 9.8
Area = π * r^2
Radius in meters = 5.4 * 10^-3 m
Area = π * (5.4 * 10^-3)^0.5

Pressure = (46 * 9.8) ÷ [π * (5.4 * 10^-3)^0.5]

Pressure = 4.921*10^6 N/m^2 (round as your instructor desires)

I hope it was helpful.. please rate me positive..